# Disturbance Compensation for Safe Kinematic Control of Robotic Systems with Closed Architecture

### v1

@@ IV-A Robot Kinematic Control with Nominal Model
The control law in (4) is ideal, assuming known dynamics and an inner-loop controller. To handle uncertainties, we propose the following practical law based on nominal dynamics:
@@
The following part shows the dynamics of the close-loop system with our nominal control law (13). We define the parametric discrepancies as:
@@
The resulting close-loop system is:
@@
Let $f$ denote the total disturbance, including internal uncertainties and external disturbances. Compared with (6), $f$ degrades system performance. To counter this, an ESO [29] is employed to estimate $f$ for disturbance rejection.
@@ IV-B Extended State Observer Design
In many robotic systems, both joint position $q$ and velocity $\dot{q}$ are measurable, enabling the design of second- or third-order ESOs based on (21). With velocity measurements available, a second-order ESO is proposed for the $i$-th joint to estimate $f_{i}$; otherwise, it can be readily extended to a third-order form.
@@
where $\hat{x}_{2i}$ and $\hat{x}_{3i}$ are the estimated velocity for $\dot{q}_{i}$ and estimated disturbance for $f_{i}$, respectively. $\beta_{1i}$ and $\beta_{2i}$ are observer gains. The estimation error dynamic is derived from (21) and (22):
@@
The observer gains are chosen such that the eigenvalues of the matrix $A_{i}$ are placed at $\omega_{o_{i}}$, where $\omega_{o_{i}}$ is the only tunable value known as the observer bandwidth [30].
@@ IV-C Controller Design
$\hat{f}$ can be used to develop a disturbance rejection controller that complements the control law (13):
@@
where $\dot{q}_{d_{0}}$ is the nominal control from (13), $\dot{q}_{d_{f}}$ rejects disturbance, and they are summed to get $\dot{q}_{d}$.
@@ IV-D Controller Stability Analysis
Assumption 3
$\left|\dfrac{\partial f_{i}(t,x,u)}{\partial t}\right|\leq l_{f}.$ (25)
Assumption 3 implies that $f$ is locally Lipschitz continuous with respect to time, and $\dot{f}$ is bounded.
@@
Given the robotic system (3) with kinematic control input, the ESO (22) with appropriate observer bandwidth $\omega_{o_{i}}$ for each joint, and the nominal control law (24) with appropriate tuned $k_{p}$ and $k_{d}$, we have the following properties: (i): the closed-loop system of (3) is bounded stable if each $\omega_{o_{i}}$ is finite; (ii): the closed-loop system of (3) is asymptotically stable if each $\omega_{o_{i}}$ approaches infinity.
@@
Since $A_{i}$ is Hurwitz, we have the following decomposition [31, Chapter 1.3]:
@@
Therefore, there exists a constant $c_{1_{i}}$ such that [31, Chapter 1.9]
@@
The matrix $H$ can be designed to be Hurwitz. Choose $V=\frac{1}{2}\epsilon^{T}Q\epsilon$ as the Lyapunov function candidate. Its derivative along (32) is given by:
$\dot{V}=\tfrac{1}{2}(\dot{\epsilon}^{T}Q\epsilon+\epsilon^{T}Q\dot{\epsilon})=-\|\epsilon\|^{2}+\xi^{T}Q\epsilon.$ (33)
@@ IV-E Robust CBF for Safe Control
By adding the disturbance term, system (11) becomes:
@@
To define safety sets for system (39), we consider an $r$-times continuously differentiable function $h(x)$, where the relative degrees of $h(x)$ with respect to both $u$ and $f$ are $r$, given that $f$ is matched. A series of functions are defined as follows:
@@
By substituting (44) into (43) and comparing it with (42), the control values in $K_{\text{cbf}}(t,x,f)$ guarantee $h_{r}(x)\geq 0$, $\forall t>0$. From the last equation in (40), we have $\dot{h}_{r-1}(x)+\gamma_{r}h_{r-1}(x)\geq 0$. Then $\dot{h}_{r-1}(x)\geq 0$ for any $x\in\partial\mathcal{C}_{r-1}$. According to Nagumo’s theorem [32], since $x(0)\in\mathcal{C}_{r-1}$, we have $h_{r-1}(x)\geq 0$, $\forall t>0$. Iteratively, $\dot{h}_{0}(x)+\gamma_{1}h_{0}(x)\geq 0$, and we have $h_{0}(x)\geq 0$, $\forall t>0$, as $x(0)\in\mathcal{C}_{0}$. Therefore, the set $\mathcal{C}_{s}$ is forward invariant for system (39). This completes the proof of Theorem 2.
@@
The estimation error of $\hat{f}_{i}$ is as follows [29]:
@@
where * represents convolution and
@@
For simplicity, let $\prod_{j=0}^{-1}(k+j)=1$. Note that (46) is the disturbance estimation error formulated in the discrete-time domain. $f_{i}(k)=f_{i}(kT_{s})$, $\hat{f}_{i}(k)=\hat{f}_{i}(kT_{s})$, $\Delta f_{i}(k)=f_{i}((k+1)T_{s})-f_{i}(kT_{s})$, $T_{s}$ is the sample time used in calculating the error bound, and $r_{i}$ and $\omega_{o_{i}}$ are the relative degree of $\dot{q}_{i}$ (output measurement) with respect to $f_{i}$ (disturbance as input) in the $i$-th joint subsystem and the poles of the $i$-th ESO in the discrete-time domain, respectively. Readers are referred to our previous work [29] for a complete proof of this error bound.
@@
As mentioned in [25], although (48) is derived in the discrete-time domain, the disturbance estimation error bound obtained from it is still the same as that from the continuous-time domain. Since we need to utilize the disturbance estimation error bound in the continuous-time domain, $T_{s}$ should be small, such as 0.1 ms, in our experiment setting. $\omega_{o_{i}}$ in the discrete-time domain needs to be converted from the continuous-time domain through Z-transform.
@@
$\begin{array}[]{r@{}l}h_{r}(x)=&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)f\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ =&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-L_{\bar{g}}L_{\psi}^{r-1}h(x)(\hat{f}-f)+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ \geq&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-|L_{\bar{g}}L_{\psi}^{r-1}h(x)|\Gamma(\omega_{o},T_{s})+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\geq 0.\end{array}$ (50)
@@
The key difference between our theoretical results in this work and our preliminary version [25] is that we have extended the robust CBF’s safety guarantee from a relative degree of one to a general case of higher order CBF.
The safety specification requires the end-effector to maintain a non-negative distance from a virtual wall. The safety function is defined as $h(x)=y-y_{0}$, where $y$ and $y_{0}$ denote the positions of the end-effector and the virtual wall, respectively. It can be verified that $h(x)$ has a relative degree of two [26]:
@@
To simultaneously achieve tracking and safe control, the nominal controller for $\dot{q}_{d}$, designed in (24), should be minimally intervened. Thus, the following QP-CBF is constructed:
@@
$+2\gamma\dot{h}(x)+\gamma^{2}h(x)\geq 0.$
V Hardware Experiments
We use the PUMA 500 robot for hardware experiments, as shown in Fig. 3. It has three main joints and a spherical wrist, providing six degrees of freedom. $q_{1}$, $q_{2}$, and $q_{3}$ correspond to the waist, shoulder, and elbow joints, while $q_{4}$, $q_{5}$, and $q_{6}$ represent the wrist rotation, bend, and flange angles. Since our focus is task-space tracking (i.e., wrist center position), only the first three joints are used. Each joint is driven by a DC motor powered by an Adept Technology MV-19 Power Chassis. The control program is built in MATLAB Simulink, which generates C code and an SDF file, and runs via a dSPACE 1103 for control and data acquisition.
@@ V Hardware Experiments
Figure 3: PUMA 500 robot used as a testbed in our research.

### v2

@@ IV Proposed Framework
This section details the controller design, along with formal proofs of stability and safety guarantee.
@@ IV-A Robot Kinematic Control with Nominal Model
The control law in (4) is ideal due to its reliance on the dynamic model and the inner-loop controller. To address uncertainties, we propose the following practical control law based on nominal values:
@@
The following part will show the resulting system with our nominal control law (15). First, we define the parametric discrepancies as:
@@
The resulting system is:
@@
We denote $f$ as the total disturbance, composed of internal dynamic uncertainty and external disturbances. Compared with (6), the presence of $f$ degrades the system performance. To mitigate this, we employ the ESO [chen2022relationship] to estimate $f$ for further disturbance rejection control.
@@ IV-B Extended State Observer Design
In many robotic systems, both the joint position $q$ and the velocity $\dot{q}$ are measurable. This allows the design of second-order and third-order ESOs based on the augmented system (23). Given the availability of velocity measurements, we propose a second-order ESO for the $i$-th joint subsystem to estimate the disturbance $f_{i}$. If velocity measurements are unavailable, the observer can be easily modified to accommodate a third-order form.
@@
where $\hat{x}_{2i}$ and $\hat{x}_{3i}$ are the estimated velocity for $\dot{q}_{i}$ and estimated disturbance for $f_{i}$, respectively. The parameters $\beta_{1i}$ and $\beta_{2i}$ are observer gains. The estimation error dynamic is derived from (23) and (24):
@@
The observer gains are chosen such that the eigenvalues of the matrix $A_{i}$ are placed at $\omega_{o_{i}}$, where $\omega_{o_{i}}$ is the only tunable value known as the observer bandwidth in the ESO design [gao2003scaling].
@@ IV-C Controller Design
The estimated disturbance $\hat{f}$ can be used to develop a disturbance rejection controller that complements the control law (15):
@@
where $\dot{q}_{d_{0}}$ is the nominal control derived from (15), $\dot{q}_{d_{f}}$ is designed to cancel the disturbance, and they are summed together to obtain the control $\dot{q}_{d}$. $\bar{M}$ and $\bar{K}_{d}$ require only nominal values, not the true values. In Sec. IV-F, we provide more details on how to estimate them.
@@ IV-D Controller Stability Analysis
Given the robotic system (3) with kinematic control input, the ESO (24) with appropriate observer bandwidth $\omega_{o_{i}}$ for each joint, and the nominal control law (26) with appropriate tuned $k_{p}$ and $k_{d}$, we have the following properties:
  1. 1.
    the closed-loop system of (3) is bounded stable if each $\omega_{o_{i}}$ is finite;
  2. 2.
    the closed-loop system of (3) is asymptotically stable if each $\omega_{o_{i}}$ approaches infinity.
@@
Since $A_{i}$ is Hurwitz, we have the following decomposition [perko2013differential, Chapter 1.3]:
@@
Therefore, there exists a constant $c_{1_{i}}$ such that [perko2013differential, Chapter 1.9]
@@
The matrix $H$ can be designed to be Hurwitz, meaning there exists a symmetric positive definite matrix $Q$ such that
$H^{T}Q+QH=-2I.$ (34)
Choose $V=\frac{1}{2}\epsilon^{T}Q\epsilon$ as the Lyapunov function candidate. Its derivative along (33) is given by:
$\begin{array}[]{ll}\dot{V}&=\frac{1}{2}(\dot{\epsilon}^{T}Q\epsilon+{\epsilon}^{T}Q\dot{\epsilon})\\ &=-\left\lVert\epsilon\right\rVert^{2}+\xi^{T}Q\epsilon.\end{array}$ (35)
@@ IV-E Robust CBF for Safe Control
By adding the disturbance term, system (13) can be written as follows:
@@
To define safety sets for system (41), we consider an $r$-times continuously differentiable function $h(x)$, where the relative degrees of $h(x)$ with respect to both $u$ and $f$ are assumed to be $r$, given that $f$ is matched. A series of functions can be defined as follows:
@@
By substituting (46) into (45) and comparing it with (44), the control values in $K_{\text{cbf}}(t,x,f)$ guarantee $h_{r}(x)\geq 0$, $\forall t>0$. From the last equation in (42), we have $\dot{h}_{r-1}(x)+\gamma_{r}h_{r-1}(x)\geq 0$. Then $\dot{h}_{r-1}(x)\geq 0$ for any $x\in\partial\mathcal{C}_{r-1}$. According to Nagumo’s theorem [blanchini2008set], since $x(0)\in\mathcal{C}_{r-1}$, we have $h_{r-1}(x)\geq 0$, $\forall t>0$. Iteratively, $\dot{h}_{0}(x)+\gamma_{1}h_{0}(x)\geq 0$, and we have $h_{0}(x)\geq 0$, $\forall t>0$, as $x(0)\in\mathcal{C}_{0}$. Therefore, the set $\mathcal{C}_{s}$ is forward invariant for system (41). This completes the proof of Theorem 2.
@@
Assumption 3
$\left|\dfrac{\partial f_{i}(t,x,u)}{\partial t}\right|\leq l_{f}.$ (48)
Assumption 3 implies that $f$ is locally Lipschitz continuous with respect to time.
The estimation error of $\hat{f}_{i}$ is as follows [chen2022relationship]:
@@
where * represents convolution and [^2] For simplicity, let $\prod_{j=0}^{-1}(k+j)=1$.
@@
Note that (49) is the disturbance estimation error formulated in the discrete-time domain. $f_{i}(k)=f_{i}(kT_{s})$, $\hat{f}_{i}(k)=\hat{f}_{i}(kT_{s})$, $\Delta f_{i}(k)=f_{i}((k+1)T_{s})-f_{i}(kT_{s})$, $T_{s}$ is the sample time used in calculating the error bound, and $r_{i}$ and $\omega_{o_{i}}$ are the relative degree of $\dot{q}_{i}$ (output measurement) with respect to $f_{i}$ (disturbance as input) in the $i$-th joint subsystem and the poles of the $i$-th ESO in the discrete-time domain, respectively. Readers are referred to our previous work [chen2022relationship] for a complete proof of this error bound, while a brief proof demonstrating that the sum of the series in (50) is bounded is provided in Appendix A.
@@
As mentioned in [chen2023robust], although (51) is derived in the discrete-time domain, the disturbance estimation error bound obtained from it is still the same as that from the continuous-time domain. Since we need to utilize the disturbance estimation error bound in the continuous-time domain, $T_{s}$ should be small, such as 0.1 ms, in our experiment setting. $\omega_{o_{i}}$ in the discrete-time domain needs to be converted from the continuous-time domain through Z-transform.
@@
$\begin{array}[]{r@{}l}h_{r}(x)=&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)f\\ &+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ =&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-L_{\bar{g}}L_{\psi}^{r-1}h(x)(\hat{f}-f)+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ \geq&L_{\psi}^{r}h(x)+L_{g}L_{\psi}^{r-1}h(x)u+L_{\bar{g}}L_{\psi}^{r-1}h(x)\hat{f}\\ &-|L_{\bar{g}}L_{\psi}^{r-1}h(x)|\Gamma(\omega_{o},T_{s})+\sum\limits_{j=0}^{r-1}k_{j}h^{(j)}(x)\\ \geq&0.\end{array}$ (53)
@@
The fundamental difference between our main theoretical results in this work and our preliminary version [chen2023robust] is that we have extended the robust CBF’s safety guarantee from a relative degree of one to a general case of higher order CBF.
Consider the safety specification for our end-effector: the relative distance between the end-effector and a virtual wall must be non-negative. The safety function can be defined as $h(x)=y-y_{0}$, where $y$ and $y_{0}$ are the positions of the end effector and the virtual wall, respectively.
Remark 3
We use this distance measurement as an example for simplification. Users can easily apply other safety specifications, such as the ball-like ones, without any fundamental difference.
It can be verified that $h(x)$ has a relative degree of two with respect to system (13) by continuously taking its derivatives. A series of functions can be defined as follows [nguyen2016exponential]:
@@
To simultaneously achieve tracking and safe control, the nominal controller for $\dot{q}_{d}$, designed in (26), should be minimally intervened. That is, the nominal control should only be overridden by the safety controller when the system is unsafe. Thus, the following QP-based controller is constructed:
@@
$+2\gamma\dot{h}(x)+\gamma^{2}h(x)\geq 0,$
IV-F Determination of Nominal Control Gain
We use the nominal model (14) throughout the experiments. A critical parameter in our controller design is the control gain in (26), $b_{0}\triangleq\bar{M}^{-1}\bar{K}_{d}$, which depends on the nominal inertia matrix $\bar{M}$ and the proportional gain of the inner-loop controller. We outline two possible approaches below for the determination of the (nominal) control gain.
[…]
Inertia matrix-based control gain tuning: Many common robotic manipulators, such as the PUMA robot used in our work, have available Universal Robot Description Format (URDF) files. A robot dynamics library can be used to parse the URDF and compute the inertia, Coriolis, and gravity matrices [carpentier2019pinocchio]. We recommend that practitioners adopt this approach to obtain the nominal inertia matrix $\bar{M}$ for the control gain $b_{0}$ without requiring system identification.
Once $\bar{M}$ is determined, tuning the control gain becomes solely a matter of adjusting $\bar{K}_{d}$. A practical rule of thumb for fine-tuning $\bar{K}_{d}$ is to start with a sufficiently large value, then gradually decrease it to improve tracking performance. A large gain generally leads to a larger but bounded tracking error while maintaining stability, whereas a small gain can reduce tracking error but increases the risk of instability if too small. The study in [xue2015performance] provides theoretical support for this trade-off. A useful tip for practitioners to simplify the fine-tuning process is to tune each joint individually rather than adjusting all coupled joints simultaneously.
In our hardware experiments, where the control input is a lower-level voltage rather than direct torque control, the control gain becomes $\bar{M}^{-1}\bar{K}_{d}B$, where $B$ represents the mapping from voltage to torque. The inertia matrix $\bar{M}$ is derived from the same parameters used in our prior work [khalaf2019trajectory], though it can also be obtained from the URDF as mentioned earlier. Even if $K_{d}$ is assumed to be known in the experiments, $B$ still requires fine-tuning. Thus, control gain tuning remains necessary in the hardware experiments. Note that in the hardware experiments, the ground-truth control gain cannot be precisely determined due to the uncertainty in $\bar{M}$ (caused by estimation errors, payload variations, etc.) and the uncertainty in $B$ (as it is a simplified linear approximation between torque and voltage). We have conducted extensive experiments to demonstrate that with a broad range of $\bar{K}_{d}$—from 0.6 to 10 times its true value—the tracking error remains within 0.2 radians, indicating that gain tuning is practical, owing to its large working range, see Sec. VI-A.
V Simulation Results
In this section, we validate our method for safety assurance in a simulation environment. The advantage of using such a fully controlled environment is that we can obtain ground truth values for all kinematic and dynamic parameters, allowing us to directly compare the safety constraints of ours with the true constraints based on the actual parameters. The PUMA 500 robot model, as shown in Fig. 3, is used for both simulation and the hardware experiments. The robot has three primary joints and a spherical wrist, providing a total of six degrees of freedom. $q_{1}$ represents the waist rotation angle, $q_{2}$ the shoulder rotation angle, and $q_{3}$ the elbow rotation angle, while $q_{4}$, $q_{5}$, and $q_{6}$ correspond to the wrist rotation angle, wrist bend angle, and flange angle, respectively. Our focus is tracking control in task space (i.e., the position of the wrist center); therefore, we only use the first three joints, as joints 4, 5, and 6 are revolute joints that do not affect the task space position.
@@ V Simulation Results
Figure 3: PUMA 500 robot used as a testbed in our research. The figure is from our previous work [khalaf2019trajectory].
The robot model and system parameters here are adopted from our previous work [khalaf2019trajectory]. The inner-loop controller is configured with a PD form, the structure of which is unknown to our outer-loop controller.
$\tau=-K_{p}({q}-{q}_{d})-K_{d}(\dot{q}-\dot{q}_{d}).$ (57)
where $K_{p}=\mathrm{diag}([1,1,1])$ and $K_{d}=\mathrm{diag}([1,1,1])$. The reference trajectory for joints $q_{1}$, $q_{2}$, and $q_{3}$ is $\begin{bmatrix}0.5\sin(t)&0.25\sin(2t)+0.5&0.25\sin(2t)+\frac{\pi}{2}\end{bmatrix}^{T}$.
In the simulation, we demonstrate how ESO compensates for disturbances to ensure safety. A virtual wall is set at $y_{0}=\begin{bmatrix}y_{1}&-0.1&y_{2}\end{bmatrix}^{T}$, allowing free movement of the end-effector along the $x$ and $z$ axes, but constraining movement along the $y$ axis to a minimum value of 0.1m-0.1m. We follow the equation in IV-E for the CBF formulation, with the parameter $\gamma=10$ in the CBF. $\omega_{o}$ is set to $80$ rad/s for each joint.
System (3) with ground truth parameters is used for simulation, neglecting the friction term, $F_{r}$ and the external disturbance, $F_{ext}$. Our nominal model is based on (14). The performance comparison is conducted using different CBFs: “True Model” refers to the CBF with the ground-truth model, “Nominal Model” refers to the CBF with the nominal model, and “Nominal Model with ESO” refers to the CBF with the nominal model plus disturbance compensation.
Figure 4: The comparison of control bounds projected into each state with the true model (blue), nominal model (red), and nominal model with ESO (yellow). ESO can narrow the gap between the nominal model and the true model.
Recall the CBF constraint in (44), which can be expressed as $u\geq LB(x)$ or $u\leq UB(x)$, where $LB(\cdot)$ and $UB(\cdot)$ represent the lower and upper bounds of the control input $u$. The determination of these bounds depends on the coefficient in front of $u$. In our case, the control input $u$ is $\dot{q}_{d}$, and the state $x$ includes the joint positions $q_{1}$, $q_{2}$, and $q_{3}$ and their derivatives. Additionally, note that the dimension of our control action is three.
Fig. 4 illustrates the bounds (lower or upper) of $\dot{q}_{d}$ based on the CBF under three scenarios: with ground truth disturbance $f$ (bounds shown as a solid blue line), robust CBF (ours, bounds shown as a dashed red line), and with only the nominal model (bounds shown as a yellow line). The y-axis represents the lower/upper bound values projected onto each state, and the x-axis represents time. As shown in the figure, the bounds generated by our approach closely match the true bounds, whereas the conventional CBF without robustness features fails to do so. The bounds play a critical role in ensuring safety for uncertain nonlinear dynamics. An inaccurate control bound in the CBF can result in overly conservative behavior or failure to ensure safety. In the hardware experiments, we will present more results with and without disturbance compensation.
VI Hardware Experiments
This section demonstrates the practical implementation of the proposed framework using the PUMA 500. Each joint of the PUMA 500 is actuated by a DC motor powered by an Adept Technology MV-19 Power Chassis. The control program is developed in MATLAB Simulink, which compiles the model and generates C code and an SDF file. Robot control and data acquisition are managed by a dSPACE 1103.
Our robotic testbed used torque-mode servo drives, meaning that motor torques are given by $\tau=Bv$, where $B$ is a diagonal gain matrix and $v$ is an analog voltage command generated by the dSPACE system. Accordingly, the system model in (1) is modified to:
${M}\ddot{q}+{C}\dot{q}+{G}+F_{r}=Bv+{J}^{T}F_{ext}.$ (58)
The inner-loop controller is configured as
$v=-K_{p}({q}-{q}_{d})-K_{d}(\dot{q}-\dot{q}_{d}).$ (59)